A FATHER IS 3 TIMES AS OLD AS HIS SON. AFTER 12 YEARS HIS AGE WILL BE TWICE AS THAT OF HIS SON THEN FIND THE PRESENT AGES

Answer by JulietG(1812)

F = 3S
F+12 = 2(S+12)
Substitute the value of F from the first equation into the second.
(3S) + 12 = 2S + 24
Subtract 2S from each side
S + 12 = 24
Subtract 12 from each side
S = 12
.
If the son is 12, then the father is 36.
In 12 years, the son will be 24 and the father 48.